# How Far Can Transformers Reason? The Globality Barrier and Inductive Scratchpad

### v2

@@ B.4 Length generalization for the parity task
Here, we provide an inductive scratchpad for the parity task which makes length generalization possible. We focus on the full parity problem, i.e., given a sequence of 00s and $1$s determining whether the number of $1$s is even (output 00) or odd (output $1$). First, we explain the input data format. We fix an embedding dimension $d_{\mathrm{emb}}$ and also assume that we have $n\leq d_{\mathrm{emb}}$ bits. Then, we choose $n$ random positions among $d_{\mathrm{emb}}$ possible ones and embed the input in dimension $d_{\mathrm{emb}}$. For the unused positions, we use a placeholder token such as a space or, here, for better presentation, an underscore. An example for $d_{\mathrm{emb}}=12$ and $n=6$ could be _01_10_0__1_.
@@
where <pointer> determines the position of the $i$th bit and <value of the $i$th bit> is the value of the $i$th bit (which can be retrieved using the pointer) and <parity of the first $i$ bits> is the cumulative parity of the first $i$ bits that can be computed using its value in the previous state (parity of the first $i-1$ bits) and the value of the $i$th bit. For the last state, state $n+1$, we use [$d_{\mathrm{emb}}$]_,<final parity><EOS> where final parity is equal to the parity of the first $n$ bits that is computed in state $n$. So, one can easily check that this scratchpad is of low globality and each state can be easily computed using the previous state and question. As an example consider input _01_10_0__1_ again. The inductive scratchpad for this example can be written as
@@
To show the length generalization ability of the inductive scratchpad we set $d_{\mathrm{emb}}=60$. Moreover, we train on samples with up to $n\leq N_{\mathrm{train}}=30$ bits. Particularly, we use fresh samples and generate each sample such that the number of bits $n$ has a uniform distribution over $\{1,2,\ldots,30\}$. We train our base decoder-only model for $2000$ iterations and we test length generalization ability for different number of bits between $N_{\mathrm{train}}=30$ and $d_{\mathrm{emb}}=60$. Results are reported in Figure 5(a). As it can be seen the model generalizes well to inputs with $50$ bits. We also observed that the generalization ability on longer sequences (e.g., $55$ bits) is not robust and depends on the random seed of the model.
@@ B.5 Length generalization for the addition task
The first format is similar to our solution for the parity length generalization problem. First, we fix an embedding dimension $d_{\mathrm{emb}}$. Assume, we want to add two operands with $n\leq d_{\mathrm{emb}}$ digits. We embed the two numbers and the plus sign $+$ in a $12d_{\mathrm{emb}}+$ 1 positions. Between them, we put spaces, or for better readability underscores. Also, we put an $=$ at the end. For example, for $x=94$, $y=31$ and $d_{\mathrm{emb}}=4$ we can have 94_+_3__1=. Now we explain the generation of the scratchpad. First, we generate a random sequence of tokens with size $d_{\mathrm{emb}}+2$ such that it begins with ‘\$’, e.g., \$xgwg6 we call this text ans[0]. Now, we enter the induction mode. For $i>0$, each state in the induction mode is given by
@@
With this scratchpad format, $N_{\mathrm{train}}=10$, and $d_{\mathrm{emb}}=30$ we can generalize to numbers with up to 18/20 digits (depending on the seed). See the results in Figure 5(b).
@@
For the second solution, we also fix an embedding dimension. $d_{\mathrm{emb}}$. Assume, we want to add two operands with $n\leq d_{\mathrm{emb}}$ digits. For input formatting, we first concatenate a \$ to the left of the two operands. Next, for each operand, we generate a random text of size $d_{\mathrm{emb}}-n$ and concatenate it to the left of the operands and call them x[0] and y[0]. Lastly, we use an $=$ to finish the question. For the scratchpad, we first generate a random text with $d_{\mathrm{emb}}+2$ tokens such that the leftmost token is \$. We call this ans[0]. Also, we use a variable for keeping the carry and we call it c which is zero at the beginning. In this format of the scratchpad, we do not use the <START> token meaning that the input question is also treated as a state (and is forgotten by the future states). We define each state ($i\geq 0$, $i=0$ corresponds to the question)
@@
where RMD represents the rightmost digit operator. In other words, at each step, we shift both the operands and the answer with the difference that operand shifts are cyclic, while for the answer we always add one digit of the correct answer to it from the left (and lose one digit from the right). Note that the position of the digits that we are adding always remains the same. Also, the addition is finished when \$ is the rightmost digit of both operands. When the addition is finished, we use one more state to output the final answer (possibly using the last carry variable). For example, for $n=2$ and $d_{\mathrm{emb}}=4$, fs\$46+ih\$98= could be an input example. In this case, the scratchpad can be written as
@@
In the example above one can note that some part of the s[0] is in the question and some part of it is in the scratchpad.
Note that the formatting required for the input of this scratchpad is stronger than the previous scratchpad. However, here we can get a much stronger length generalization. By setting $d_{\mathrm{emb}}=30$ and training on numbers with up to $4$ digits, we can generalize to numbers with up to $26$ digits and even $30$ digits for some of the seeds (see Figure 5(b)).
@@ C.2 Implementation of the inductive scratchpad
where we have assumed that $t$ tokens have appeared before the first state s[1]. Note that this number may vary across samples. We can easily create an attention mask based on the attention groups. All groups only attend to tokens in their group and tokens in group 00 (everything that comes before <START>). Also, using the loss mask vector, we do not compute the loss for the second copy of each state. Also, for each state generation, we reset the positional indices of the tokens. Using the appropriate loss mask, attention mask, and positional indices explained above guarantees that the loss is computed exactly once for each state, and each state is generated based on the previous state and tokens before <START> (i.e., the question Q) in an identical manner. Note that both approaches that we discussed for the train time can be implemented easily for conventional Transformer models. Further, they do not change Transformer models’ behavior on non-inductive samples. They also work with both absolute and relative positional embeddings. Note that the first approach favors Transformers with a small block size (context length) and the second approach is more suitable when the block size is large. So based on the block size, a mixture of these two approaches can be used. In our implementation, we use the second approach as we do not have an issue with the block size.
@@ C.3 Comparison with relative positional embeddings
Appendix D Proof of Theorem 1
@@
The positional encodings are iid and there is no causal masking, so the probability distribution of the Transformer is symmetric under permutations of the inputs. In particular, for an arbitrary permutation $\pi\in S_{n}$ we can move all of the letters in block $i$ to the corresponding position in block $\pi(i)$ for each $i$ to essentially apply this permutation to the $\sigma$. Now, let $Y$ be the indicator function for $\prod_{i=1}^{n}\sigma_{i}$ being the identity. Then, for every select $z\in\{0,1\}^{\lfloor(n-1)/2\rfloor}$ let $\pi_{z}=\prod_{i\leq\lfloor(n-1)/2\rfloor:z_{i}=1}(2i-1,2i)$ and $Y_{z}$ be the indicator function for $\prod_{i=1}^{n}\sigma_{\pi_{z}(i)}$ being the identity. By the symmetries of the Transformer, it will have as hard a time learning to compute $Y$ as it would have learning to compute $Y_{z}$ for any $z$. Next, let $Y_{\emptyset}$ be $1$ with probability $1/3$ and 00 with probability $2/3$, independently of $X$. We claim that the probability distribution of the Transformer’s weights when it is trained on $Y_{z}$ will be essentially the same as its weights when it is trained on $Y_{\emptyset}$. In order to formalize that, let $T_{w}(x)$ be the output the Transformer gives when its edge weights and positional encoding take on the values given by $w$ on an input of $x$. Next, let $P_{W,\emptyset,t}$ be the probability distribution of the weights and positional embeddings after $t$ time steps of training on $Y_{\emptyset}$ and $P_{W,z,t}$ be the probability distribution of the weights and positional embeddings after $t$ time steps of training on $Y_{z}$. For each $t$, $P_{W,\emptyset,t}$ is symmetric under permutations of the positional embeddings.
@@
$=\mathbb{E}_{w\sim P_{W,\emptyset,t}}[\mathbb{E}_{X,Y_{\emptyset}}[L^{\prime}(Y_{\emptyset},T_{w}(X))]]+\mathbb{E}_{w\sim P_{W,\emptyset,t}}[\mathbb{E}_{z}[\mathbb{E}_{X}[(Y-1/3)\cdot(L^{\prime}(1,T_{w}(X))-L^{\prime}(0,T_{w}(X)))]]]$
$\geq\mathbb{E}_{w\sim P_{W,\emptyset,t}}[\mathbb{E}_{X,Y_{\emptyset}}[L^{\prime}(Y_{\emptyset},y_{0})]]-\mathbb{E}_{w\sim P_{W,\emptyset,t}}\left[\sqrt{\mathbb{E}_{z}[\mathbb{E}^{2}_{X}[(Y-1/3)\cdot(L^{\prime}(1,T_{w}(X))-L^{\prime}(0,T_{w}(X)))]]}\right]$
@@
D.1 Extension to agnostic scratchpads
Theorem 1 can also be generalised to Transformers trained with agnostic scratchpads in order to get the following.
@@ D.1 Extension to agnostic scratchpads
The positional encodings are iid and there is no causal masking of the original input, so the probability distribution of the Transformer is symmetric under permutations of the inputs. In particular, for an arbitrary permutation $\pi\in S_{n}$ we can move all of the letters in block $i$ to the corresponding position in block $\pi(i)$ for each $i$ to essentially apply this permutation to the $\sigma$. Now, let $Y$ be the indicator function for $\prod_{i=1}^{n}\sigma_{i}$ being the identity. Then, for every $z\in\{0,1\}^{\lfloor(n-1)/2\rfloor}$ let $\pi_{z}=\prod_{i\leq\lfloor(n-1)/2\rfloor:z_{i}=1}(2i-1,2i)$ and $Y_{z}$ be the indicator function for $\prod_{i=1}^{n}\sigma_{\pi_{z}(i)}$ being the identity. By the symmetries of the Transformer, it will have as hard a time learning to compute $Y$ as it would have learning to compute $Y_{z}$ for any $z$. Next, let $Y_{\emptyset}$ be $1$ with probability $1/3$ and 00 with probability $2/3$, independently of $X$. We claim that the probability distribution of the Transformer’s weights when it is trained on $Y_{z}$ will be essentially the same as its weights when it is trained on $Y_{\emptyset}$. In order to formalize that, let $T_{w}(x)$ be a random output the Transformer with scratchpad gives when its edge weights and positional encoding take on the values given by $w$ on an input of $x$. Next, let $P_{W,\emptyset,t}$ be the probability distribution of the weights and positional embeddings after $t$ time steps of training on $Y_{\emptyset}$ and $P_{W,z,t}$ be the probability distribution of the weights and positional embeddings after $t$ time steps of training on $Y_{z}$. For each $t$, $P_{W,\emptyset,t}$ is symmetric under permutations of the positional embeddings.
@@
Appendix E Comment on Lemma 1

### v3

@@ B.4 Length generalization for the parity task
Here, we provide an inductive scratchpad for the parity task which makes length generalization possible. We focus on the full parity problem, i.e., given a sequence of 00s and $1$s determining whether the number of $1$s is even (output 00) or odd (output $1$).
First, we explain the input data format. We fix an ambient dimension $d_{\mathrm{amb}}$ and also assume that we have $n\leq d_{\mathrm{amb}}$ bits. Then, we choose $n$ random positions among $d_{\mathrm{amb}}$ possible ones and embed the input in dimension $d_{\mathrm{amb}}$. For the unused positions, we use a placeholder token such as a space or, here, for better presentation, an underscore. An example for $d_{\mathrm{amb}}=12$ and $n=6$ could be _01_10_0__1_.
@@
where <pointer> determines the position of the $i$th bit and <value of the $i$th bit> is the value of the $i$th bit (which the model can learn to retrieve using the pointer) and <parity of the first $i$ bits> is the cumulative parity of the first $i$ bits that can be computed using its value in the previous state (parity of the first $i-1$ bits) and the value of the $i$th bit. For the last state, state $n+1$, we use [$d_{\mathrm{amb}}$]_,<final parity><EOS> where final parity is equal to the parity of the first $n$ bits that is computed in state $n$. So, one can easily check that this scratchpad is of low globality and each state can be easily computed using the previous state and question. As an example consider input _01_10_0__1_ again. The inductive scratchpad for this example can be written as
@@
where different states are shown by different colors. To show the length generalization ability of the inductive scratchpad we set $d_{\mathrm{amb}}=60$. Moreover, we train on samples with up to $n\leq N_{\mathrm{train}}=30$ bits. Particularly, we use fresh samples and generate each sample such that the number of bits $n$ has a uniform distribution over $\{1,2,\ldots,30\}$. We train our base decoder-only model for $2000$ iterations and we test length generalization ability for different number of bits between $N_{\mathrm{train}}=30$ and $d_{\mathrm{amb}}=60$. Results are reported in Figure 5(a). As it can be seen the model generalizes well to inputs with $50$ bits. We also observed that the generalization ability on longer sequences (e.g., $55$ bits) is not robust and depends on the random seed of the model.
@@ B.5 Length generalization for the addition task
The first format is similar to our solution for the parity length generalization problem. First, we fix an ambient dimension $d_{\mathrm{amb}}$. Assume, we want to add two operands with $n\leq d_{\mathrm{amb}}$ digits. We embed the two numbers and the plus sign $+$ in a $12d_{\mathrm{amb}}+$ 1 positions. Between them, we put spaces, or for better readability underscores. Also, we put an $=$ at the end. For example, for $x=94$, $y=31$ and $d_{\mathrm{amb}}=4$ we can have 94_+_3__1=. Now we explain the generation of the scratchpad. First, we generate a random sequence of tokens with size $d_{\mathrm{amb}}+2$ such that it begins with ‘\$’, e.g., \$xgwg6 we call this text ans[0]. Now, we enter the induction mode. For $i>0$, each state in the induction mode is given by
@@
where different states are represented in different colors. With this scratchpad format, $N_{\mathrm{train}}=10$, and $d_{\mathrm{amb}}=30$ we can generalize to numbers with up to 18/20 digits (depending on the seed). See the results in Figure 5(b).
@@
For the second solution, we also fix an ambient dimension. $d_{\mathrm{amb}}$. Assume, we want to add two operands with $n\leq d_{\mathrm{amb}}$ digits. For input formatting, we first concatenate a \$ to the left of the two operands. Next, for each operand, we generate a random text of size $d_{\mathrm{amb}}-n$ and concatenate it to the left of the operands and call them x[0] and y[0]. Lastly, we use an $=$ to finish the question. For the scratchpad, we first generate a random text with $d_{\mathrm{amb}}+2$ tokens such that the leftmost token is \$. We call this ans[0]. Also, we use a variable for keeping the carry and we call it c which is zero at the beginning. In this format of the scratchpad, we do not use the <START> token meaning that the input question is also treated as a state (and is forgotten by the future states). We define each state ($i\geq 0$, $i=0$ corresponds to the question)
@@
where RMD represents the rightmost digit operator. In other words, at each step, we shift both the operands and the answer with the difference that operand shifts are cyclic, while for the answer we always add one digit of the correct answer to it from the left (and lose one digit from the right). Note that the position of the digits that we are adding always remains the same. Also, the addition is finished when \$ is the rightmost digit of both operands. When the addition is finished, we use one more state to output the final answer (possibly using the last carry variable). For example, for $n=2$ and $d_{\mathrm{amb}}=4$, fs\$46+ih\$98= could be an input example. In this case, the scratchpad can be written as
@@
where different colors are used for different states. In the example above one can note that some part of the s[0] is in the question and some part of it is in the scratchpad.
Note that the formatting required for the input of this scratchpad is stronger than the previous scratchpad. However, here we can get a much stronger length generalization. By setting $d_{\mathrm{amb}}=30$ and training on numbers with up to $4$ digits, we can generalize to numbers with up to $26$ digits and even $30$ digits for some of the seeds (see Figure 5(b)).
@@ C.2 Implementation of the inductive scratchpad
where we have assumed that $t$ tokens have appeared before the first state s[1]. Note that this number may vary across samples. We can easily create an attention mask based on the attention groups. All groups only attend to tokens in their group and tokens in group 00 (everything that comes before <START>). Also, using the loss mask vector, we do not compute the loss for the second copy of each state (where the loss mask is equal to zero). Also, for each state generation, we reset the positional indices of the tokens. Using the appropriate loss mask, attention mask, and positional indices explained above guarantees that the loss is computed exactly once for each state, and each state is generated based on the previous state and tokens before <START> (i.e., the question Q) in an identical manner. Note that both approaches that we discussed for the train time can be implemented easily for conventional Transformer models. Further, they do not change Transformer models’ behavior on non-inductive samples. Our methods also work with both absolute and relative positional embeddings. Note that the first approach favors Transformers with a small block size (context length) and the second approach is more suitable when the block size is large. So based on the block size, a mixture of these two approaches can be used. In our implementation, we use the second approach as we do not have an issue with the block size.
@@
We note that in the inductive scratchpad, all the states except the penultimate one are ignored. As a result, the effective number of tokens is significantly reduced compared to the full scratchpad. Consequently, the inductive scratchpad works better for longer scratchpads and models with more limited context sizes and generally scales better.
Also, note that the inductive scratchpad can generally be used for a wide range of algorithmic tasks. For example, consider an algorithm with a for loop that updates some variables. One can easily put the variables of the algorithm in the state of an inductive scratchpad, and use the Transformer for computing the values at each iteration (and also determining the halting of the loop). The inductive scratchpad for the parity, addition, and cycle tasks all fall into this category.
@@ C.3 Comparison with relative positional embeddings
Appendix D Further discussion and specification of Conjecture 1
Definition 6 (Well-behaved distribution for Conjecture 1).
Input distribution $P_{X}$ over alphabet $\mathcal{A}^{n}$ with $|\mathcal{A}|=O(n^{c})$ is well-behaved if for $X\sim P_{X}$ there is no value that $X$ takes with probability $\Omega(n^{-c_{1}})$ for any $c_{1}>0$, and every eigenvalue of $X$’s covariance matrix of the indicator functions for the values of $X$’s entries has absolute value $O(n^{-c_{2}}\sum_{i=1}^{n}\mathrm{Var}(X_{i}))$ for some $c_{2}>0$.
In other words, the first requirement means that there is no value of $X$ that is frequent enough that allows the model to weakly learn the function simply by memorizing the value of that input. The second rules out the possibility of an input distribution with a few important components that are easy to notice and determine the output, such as a scenario where there are $\log(n)$ blocks of $n/(2\log(n))$ bits such that all of the bits in a block are always the same and the output depends only on the values of the blocks. Other than that, many distributions of interest are well-behaved, such as the i.i.d. measure as in [12, 25, 24, 28, 26, 27], or measures with dependencies as for the cycle task considered in this paper.
Intuition on Conjecture 1. To see the role of the histogram, $\hat{P}_{X}$, note that if one removes positional embeddings from a Transformer, the Transformer would become permutation invariant. In this case, learning functions like full parity (parity of all the bits) becomes very easy as there are only $n+1$ possibilities for $n$ bits in the eyes of a permutation invariant model. One could potentially achieve a similar effect to removing the positional embeddings, if one initializes positional embeddings with a small enough (or vanishing) scale. That is the reason that $\hat{P}_{X}$ needs to be included in the definition of the globality degree. Further see Appendix F for more intuition on Conjecture 1.
Appendix E Intuition on the agnostic scratchpad
In a lot of the cases where there is a scratchpad, we assume that during training we know what the Transformer should write in the scratchpad and just need to teach it to do so. However, this tends to be an unrealistic assumption in practice. If we knew how to compute the appropriate scratchpad entries from the inputs then we could just give the Transformer a prefilled scratchpad instead of training it to write the appropriate entries in the scratchpad. Admittedly, we could have a scenario where it is expensive for us to compute the appropriate scratchpad entries and we are hoping to teach the Transformer to do it for us, but there is a limited middle ground between it being cheap enough that we do not need the Transformer to compute the entries and it being too expensive to compute the correct scratchpad entries for all the training data. The other case where we would know what the correct scratchpad entries were during training is if the training data already came with scratchpad entries, which could happen if our training data includes some kind of explanation, which might not often be the case.
So, we could easily have a case where we are trying to train a Transformer with a scratchpad but do not know what it should write in the scratchpad. In that case, it seems like the best available way to evaluate whether an entry the model wrote in the scratchpad is right or not is to judge it based on the end result of writing that entry. In other words, we define a loss function for entries in the scratchpad by taking the scratchpad with that entry, extending it to a full scratchpad, and checking if the transformer ends up giving the right answer. This is the mechanism behind the algorithm proposed in Conjecture 2 for optimizing the network based on agnostic scratchpads.
Appendix F Proof of Theorem 1
@@
The positional encodings are iid and there is no causal masking, so the probability distribution of the Transformer is symmetric under permutations of the inputs. In particular, for an arbitrary permutation $\pi\in S_{n}$ we can move all of the letters in block $i$ to the corresponding position in block $\pi(i)$ for each $i$ to essentially apply this permutation to the $\sigma$. Now, let $Y$ be the indicator function for $\prod_{i=1}^{n}\sigma_{i}$ being the identity. Then, for every select $z\in\{0,1\}^{\lfloor(n-1)/2\rfloor}$ let $\pi_{z}=\prod_{i\leq\lfloor(n-1)/2\rfloor:z_{i}=1}(2i-1,2i)$ and $Y_{z}$ be the indicator function for $\prod_{i=1}^{n}\sigma_{\pi_{z}(i)}$ being the identity. By the symmetries of the Transformer, it will have as hard a time learning to compute $Y$ as it would have learned to compute $Y_{z}$ for any $z$. Next, let $Y_{\emptyset}$ be $1$ with probability $1/3$ and 00 with probability $2/3$, independently of $X$. We claim that the probability distribution of the Transformer’s weights when it is trained on $Y_{z}$ will be essentially the same as its weights when it is trained on $Y_{\emptyset}$. In order to formalize that, let $T_{w}(x)$ be the output the Transformer gives when its edge weights and positional encoding take on the values given by $w$ on an input of $x$. Next, let $P_{W,\emptyset,t}$ be the probability distribution of the weights and positional embeddings after $t$ time steps of training on $Y_{\emptyset}$ and $P_{W,z,t}$ be the probability distribution of the weights and positional embeddings after $t$ time steps of training on $Y_{z}$. For each $t$, $P_{W,\emptyset,t}$ is symmetric under permutations of the positional embeddings.
@@
$=\mathbb{E}_{w\sim P_{W,\emptyset,t}}[\mathbb{E}_{X,Y_{\emptyset}}[L^{\prime}(Y_{\emptyset},T_{w}(X))]]$
$\geq\mathbb{E}_{w\sim P_{W,\emptyset,t}}[\mathbb{E}_{X,Y_{\emptyset}}[L^{\prime}(Y_{\emptyset},y_{0})]]$
$\phantom{xxxxxx}-\mathbb{E}_{w\sim P_{W,\emptyset,t}}\left[\sqrt{\mathbb{E}_{z}[\mathbb{E}^{2}_{X}[(Y-1/3)\cdot(L^{\prime}(1,T_{w}(X))-L^{\prime}(0,T_{w}(X)))]]}\right]$
@@
F.1 Extension to agnostic scratchpads
Theorem 1 can also be generalized to Transformers trained with agnostic scratchpads in order to get the following.
@@ F.1 Extension to agnostic scratchpads
The positional encodings are iid and there is no causal masking of the original input, so the probability distribution of the Transformer is symmetric under permutations of the inputs. In particular, for an arbitrary permutation $\pi\in S_{n}$ we can move all of the letters in block $i$ to the corresponding position in block $\pi(i)$ for each $i$ to essentially apply this permutation to the $\sigma$. Now, let $Y$ be the indicator function for $\prod_{i=1}^{n}\sigma_{i}$ being the identity. Then, for every $z\in\{0,1\}^{\lfloor(n-1)/2\rfloor}$ let $\pi_{z}=\prod_{i\leq\lfloor(n-1)/2\rfloor:z_{i}=1}(2i-1,2i)$ and $Y_{z}$ be the indicator function for $\prod_{i=1}^{n}\sigma_{\pi_{z}(i)}$ being the identity. By the symmetries of the Transformer, it will have as hard a time learning to compute $Y$ as it would have to learn to compute $Y_{z}$ for any $z$. Next, let $Y_{\emptyset}$ be $1$ with probability $1/3$ and 00 with probability $2/3$, independently of $X$. We claim that the probability distribution of the Transformer’s weights when it is trained on $Y_{z}$ will be essentially the same as its weights when it is trained on $Y_{\emptyset}$. In order to formalize that, let $T_{w}(x)$ be a random output the Transformer with scratchpad gives when its edge weights and positional encoding take on the values given by $w$ on an input of $x$. Next, let $P_{W,\emptyset,t}$ be the probability distribution of the weights and positional embeddings after $t$ time steps of training on $Y_{\emptyset}$ and $P_{W,z,t}$ be the probability distribution of the weights and positional embeddings after $t$ time steps of training on $Y_{z}$. For each $t$, $P_{W,\emptyset,t}$ is symmetric under permutations of the positional embeddings.
@@
Appendix G Comment on Lemma 1